Determine the value of the unknown in the given matrix equalities



(2 −1 x


4 7 0)


(x 3


−2 5


1 −4)


=


(4 −2


1 5 )


1
Expert's answer
2022-06-12T16:55:25-0400
"\begin{pmatrix}
 2 & -1 & x \\
 4 & 7 & 0
\end{pmatrix}\begin{pmatrix}
 x & 3 \\
 -2 & 5 \\
1 & -4
\end{pmatrix}"

"=\begin{pmatrix}
 2x+2+x & 6-5-4x \\
 4x-14 & 12+35
\end{pmatrix}"

"=\begin{pmatrix}
 3x+2 & 1-4x \\
 4x-14 & 47
\end{pmatrix}"


"\begin{pmatrix}
 3x+2 & 1-4x \\
 4x-14 & 47
\end{pmatrix}\not=\begin{pmatrix}
 4 & -2 \\
 1 & 5
\end{pmatrix}, x\in \R"

There is no solution.